OpenStudy (anonymous):

trig identity help!!! Solve the following quadratic trig equations algebraically (without graphing technology) for the domain 0º < x < 360º. Use exact solutions whenever possible. a) 6cos2( theta) + cos(theta ) - 1 = 0 b) 2 - 2sin2(theta ) = cos( theta)

OpenStudy (amistre64):

6cos(2x) + cos(x) -1 = 0? or 6cos(x)^2 + cos(x) -1 =0? ill assume this one. (6)(-1) = -6 factors of (-6) that equal (+1): (3)(-2) (cos(x)+(3/6)) (cos(x)-(2/6)) cos(x) =(-1/2) or cos(x)=(1/3) x = 120 and 240 for cos(-1/2) and x= some angle in Q1 and Q2 that has a cos^-1(1/3) if I did it right :)